# Stochastic Reachability of Uncontrolled Systems via Probability Measures: Approximation via Deep Neural Networks

### v2

@@ I Introduction
We propose an alternative framework for stochastic reachability based in probability measures. The probability measure of a random variable returns, for a set input (e.g., a set of states), the likelihood of the variable lying in that set. With a probability measure, we can find a set associated with any prescribed likelihood [10, 11]. Occupation measures, a type of probability measure, have been used to compute regions of attraction [12] and control invariant sets [13]. Here, we propose a measure theoretic formalism of the stochastic reachability problem to represent the probability measures of the state. We focus on autonomous systems, that is, dynamical systems with no control input. One of the main challenges in constructing this formalism is the need for assurances of constraint satisfaction.
The main contribution of this paper is a formulation of the stochastic reachability problem via probability measures, that captures the state distribution which satisfies the reachability specification at every time step. We synthesize a probability measure that ensures compliance with reachability specifications for all probabilities. We additionally seek numerical implementations that can approximate stochastic reachable sets by exploiting this formalism. We capture the level sets for some likelihood as a root-finding problem over possible probability measures, then propose a data-driven approach based in empirical assessments of constraint satisfaction. This approach reduces problem of constructing a state probability measure to one of forward propagation of the dynamics, to empirically evaluate probabilities.
Lastly, we employ a deep neural net to learn the level sets of the state probability measure for a finite time horizon, so that given a particular state and time step to evaluate, the neural net predicts the stochastic reachability probability. Our deep neural net learns the probability measure across all safety likelihoods simultaneously, but has computational complexity that scales with the network size (as determined by the number of neurons and layers).
@@ II Preliminaries and Problem Formulation
${1}_{T}(x_{k+1})=\begin{cases}1,\ x\in T\\ 0,\ x\notin T.\end{cases}$ (2)
@@
We note that $\mu_{k+1}$ is unique [14, Lemma 10.4.3].
@@ III-A Backward Recursion of the State Probability Measure
  • (5) holds for $k\in\mathbb{N}_{[0,N-2]}$,
  • (6) holds for $k=N-1$,
@@ IV-A Forming the Training Data
Here, $L$ denotes the number of samples of the subsequent states $x_{j,k+1}$ starting from a state $x_{k}$ that is sampled from a uniform distribution over the state space, i.e. $x_{k}\sim\mathrm{Uniform}(\mathcal{X})$. We restrict our attention to a hyperrectangle in the state space, taking $\mathcal{X}=[-b,b]^{n}$ with bounds $b\in\mathbb{R}$, which we assume encompass the tube sets $T_{k}$. This process is contingent on the fact that since we are handling dynamical systems which have no control input, we can avoid a backward recursion via dynamic programming and, instead, employ forward simulation to obtain an empirical estimate. We also note that this process would not be effective for a dynamical system with a control input.
@@
Let (9) denote the true value function that is defined via a Lebesgue integral [14, Ch. 2, Definition 2.4.1] of a function with respect to the previous transition kernels as well as prior state probability measures and let (14) denote an empirical value function averaged over $L$ samples detailed in Algorithm 1. The strong law of large numbers [14, Ch. 10, 10.10(v)] ensures, as $L$ increases, $\hat{V}_{k}$ will converge almost everywhere to $V_{k}$. Thus, for $k\in\mathbb{N}_{[0,N-1]}$, $\lim_{L\rightarrow\infty}\hat{V}_{k}(x_{k})=V_{k}(x_{k})$. ∎
@@
Figure 2: We compare our approach (left) to the ground truth (middle) via dynamic programming (DP) as well as to a reproducing kernel Hilbert space (RKHS) approach (right)  [9]. In the top row, for a Gaussian disturbance with $\sigma^{2}=0.01$, the RKHS outperforms our method, matching closely with DP, as the neural network struggles to learn the sharp drop off of probabilities at the boundaries. However, in the second row, when the Gaussian disturbance has $\sigma^{2}=0.1$ variance, our approach is close to the DP solution in comparison to RKHS. Note that we are able to train the neural network with more samples ($M=5000$ state samples at training time with $L=2000$ to compute $\alpha$ offline) versus the RKHS method ($M=10,201$ samples at training time). When attempting to use the same number of samples ($M\cdot L=10,000,000$ samples) for the RKHS approach, we ran out of memory.
@@ IV-B Using DNNs as the Function Approximator
Neural Network Structure: We design our DNN as a fully-connected, feed-forward neural network.
We use $4$ hidden layers where each layer consists of $64$ neurons, each employing the Rectified Linear Unit (ReLU) activation function. The input into the neural network is the state, $x\in\mathcal{X}$, and the current time step, $k\in\mathbb{N}$.
Before turning to the examples, we briefly consider the matter of computational complexity involved in evaluating the function approximator. Note that DNNs evaluations scale with the architecture’s size, specifically the arrangement and connectivity of its layers, and not the sample size. This means that a trained neural network can accommodate additional data without a proportional increase in the evaluation time. In comparison, kernel methods, which scale with the number of samples, struggle at evaluation time due to the size of the Gram matrix [15]. Although kernel methods have seen scalability improvements via random Fourier features [16], neural networks have made parallel strides in scalability through techniques such as quantization [17], to represent neural networks on a smaller memory footprint, and batching of the data for training with limited memory [18, Ch. 8].
@@ VI Conclusion
In this paper, we provide a measure theoretic formalism of the stochastic reachability problem for uncontrolled systems, that characterizes stochastic reachability in terms of probability measures. This approach enables new computational tools for approximate solutions to stochastic reachable sets, by computing level sets of the probability measure for all probability values. We develop a numerical implementation that employs deep neural nets to approximate the stochastic reachability probability, for a given state and time step within a finite time horizon. Future work will focus on extensions to accommodate dynamical systems with control inputs.

### v3

@@ I Introduction
The main contribution of this paper is a formulation of the stochastic reachability problem via probability measures, that captures the state distribution which satisfies the reachability specification at every time step. We synthesize a probability measure that ensures compliance with reachability specifications for all probabilities. We propose a measure-based approach to approximately computing reachable sets using neural nets as surrogate functions. The probability measure of a random variable returns, for a set input (e.g., a set of states), the likelihood of the variable lying in that set. With a probability measure, we can find a set associated with any prescribed likelihood [10, 11]. Occupation measures, a type of probability measure, have been used to compute regions of attraction [12] and control invariant sets [13]. Essentially, we use measures to represent the stochastic nature of the evolution of the system state. We additionally seek numerical implementations that can approximate stochastic reachable sets by exploiting this formalism. We capture the level sets for some likelihood as a root-finding problem over possible probability measures, then propose a data-driven approach based in empirical assessments of constraint satisfaction. This approach reduces problem of constructing a state probability measure to one of forward propagation of the dynamics, to empirically evaluate probabilities. We construct these measures using a backwards recursion formalism; this formalism yields a sequence of value functions that we estimate from forwards trajectory data using a neural network surrogate function. The value functions themselves are not novel– indeed, they are central to the dynamic programming approach of [1]– but they are intensive to compute in general. The method by which we approximate them from data, and to use them to generalize the observed evolutions into an approximate reachable set over the continuum of the state space, is another of our contributions. This approach to approximate reachability has proven successful in other domains where dynamic programming is employed, e.g. in [14], but has not yet been applied to the case of stochastic reachability. We focus on autonomous systems, that is, dynamical systems with no control input.
Our deep neural net learns the probability measure across all safety likelihoods simultaneously, but has computational complexity that scales with the network size (as determined by the number of neurons and layers).
@@ II Preliminaries and Problem Formulation
${1}_{T}(x)=\begin{cases}1,\ x\in T\\ 0,\ x\notin T.\end{cases}$ (2)
@@
We note that $\mu_{k+1}$ is unique [15, Lemma 10.4.3].
@@ III-A Backward Recursion of the State Probability Measure
  • (5) holds for $k\in\mathbb{N}_{[0,N-2]}$;
  • (6) holds;
@@ IV-A Forming the Training Data
Here, $L$ denotes the number of samples of the subsequent states $x_{k+1,j}$ starting from a state $x_{k}$ that is sampled from a uniform distribution over the state space, i.e. $x_{k}\sim\mathrm{Uniform}(\mathcal{X})$. We restrict our attention to a hyperrectangle in the state space, taking $\mathcal{X}=[-b,b]^{n}$ with bounds $b\in\mathbb{R}$, which we assume encompass the tube sets $T_{k}$. This process is contingent on the fact that since we are handling dynamical systems which have no control input, we can avoid a backward recursion via dynamic programming and, instead, employ forward simulation to obtain an empirical estimate. We also note that this process would not be effective for a dynamical system with a control input.
@@
Let (9) denote the true value function that is defined via a Lebesgue integral [15, Ch. 2, Definition 2.4.1] of a function with respect to the previous transition kernels as well as prior state probability measures and let (14) denote an empirical value function averaged over $L$ samples detailed in Algorithm 1. The strong law of large numbers [15, Ch. 10, 10.10(v)] ensures, as $L$ increases, $\hat{V}_{k}$ will converge almost everywhere to $V_{k}$. Thus, for $k\in\mathbb{N}_{[0,N-1]}$, $\lim_{L\rightarrow\infty}\hat{V}_{k}(x_{k})=V_{k}(x_{k})$. ∎
@@
Figure 2: We compare our approach to the ground truth via dynamic programming (DP) as well as to a reproducing kernel Hilbert space (RKHS) approach [9]. In the top row, for a Gaussian disturbance with $\sigma^{2}=0.01$, the RKHS (right) outperforms our method (left), matching closely with DP (middle), as the neural network struggles to learn the sharp drop off of probabilities at the boundaries. However, in the second row, when the Gaussian disturbance has $\sigma^{2}=0.1$ variance, our approach (left) is close to the DP solution (middle) in comparison to RKHS (right). Note that we are able to train the neural network with more samples ($M=5000$ state samples at training time with $L=2000$ to compute $\alpha$ offline) versus the RKHS method ($M=10,201$ samples at training time). When attempting to use the same number of samples ($M\cdot L=10,000,000$ samples) for the RKHS approach, we ran out of memory.
@@ IV-B Using DNNs as the Function Approximator
Neural Network Structure: We design our DNN as a fully-connected, feed-forward neural network. We use $4$ hidden layers where each layer consists of $64$ neurons, each employing the Rectified Linear Unit (ReLU) activation function. The input into the neural network is the state, $x\in\mathcal{X}$, and the current time step, $k\in\mathbb{N}$.
Before turning to the examples, we briefly consider the matter of computational complexity involved in evaluating the function approximator. Note that DNNs evaluations scale with the architecture’s size, specifically the arrangement and connectivity of its layers, and not the sample size. This means that a trained neural network can accommodate additional data without a proportional increase in the evaluation time. In comparison, kernel methods, which scale with the number of samples, struggle at evaluation time due to the size of the Gram matrix [16]. Although kernel methods have seen scalability improvements via random Fourier features [17], neural networks have made parallel strides in scalability through techniques such as quantization [18], to represent neural networks on a smaller memory footprint, and batching of the data for training with limited memory [19, Ch. 8].
@@ V-B Quaternion Attitude Dynamics
The differences in runtime between the RKHS and DNN approaches in the double integrator and attitude dynamics examples are influenced by the complexity of the systems and number of samples used. In the double integrator example, the RKHS takes longer to compute due to the cost of evaluating a large Gram matrix with more samples. However, in the attitude dynamics example, the RKHS method is quicker, likely because the kernel evaluations and matrix operations are less computationally intensive for this particular sample size of 2500 compared to the DNN’s extensive training requirements.
@@ VI Conclusion
We provided a measure theoretic formalism of the stochastic reachability problem for uncontrolled systems, that characterizes stochastic reachability in terms of probability measures. This approach enables new computational tools for approximate solutions to stochastic reachable sets, by computing level sets of the probability measure for all probability values. We developed a numerical implementation that employs deep neural nets to approximate the stochastic reachability probability, for a given state and time step within a finite time horizon. Future work will focus on extensions to accommodate dynamical systems with control inputs. The essence of this extension will be to generalized the transition kernel framing of the dynamics into one that admits inputs, and to extend the sampling procedure to sample over inputs as well. This will require a restriction of the space of control inputs, at least initially, to a finite dimensional space, e.g. to piecewise constant signals.